OpenStudy (anonymous):

Find the derivative of f(x) = 3x + 8 at x = 4.

terenzreignz (terenzreignz):

Interestingly (though not really)... the derivative of a linear function is just the slope.

OpenStudy (inkyvoyd):

It's delta y/delta x as delta y and delta x get close to zero - but for a linear equation, the slope is always constant - aka the rate of change

OpenStudy (anonymous):

these are the answers idk how to get it.. 1 3 4 8

OpenStudy (inkyvoyd):

word of unrelated advice - learn calculus. Seriously, if you're in AB or BC and you decide for whatever reason not to pay attention during limits and derivatives, you're going to fail integrals completely

terenzreignz (terenzreignz):

<le gasp> First things first... the derivative of a sum (or difference) is the *sum (difference) of their derivatives* Aye?

OpenStudy (anonymous):

ya seriously lol i'm in pre calc..... and i think so

terenzreignz (terenzreignz):

Okay, what's the derivative of a power? \[\Large \frac{d}{dx}ax^n=\color{red}?\]

OpenStudy (anonymous):

0?

terenzreignz (terenzreignz):

Don't guess. You seriously aren't familiar with the power rule? D:

OpenStudy (anonymous):

nope i'm taking it online they don't teach anything

terenzreignz (terenzreignz):

Here it is... \[\Large \frac{d}{dx}ax^n\] You take the exponent (n), multiply it to the function: \[\Large \frac{d}{dx}ax^n\rightarrow \color{red}nax^n\] BUT you subtract 1 from the exponent \[\Large \frac{d}{dx}ax^n\rightarrow \color{red}nax^n\rightarrow nax^{\color{red}{n-1}}\]

terenzreignz (terenzreignz):

And THAT... is the power rule: \[\Large \frac{d}{dx}ax^n = \color{blue}{nax^{n-1}}\]

terenzreignz (terenzreignz):

So, having said that, what is the derivative of \[\Large \frac{d}{dx}3x = \frac{d}{dx}3x^1=\color{red}?\]

OpenStudy (anonymous):

is it 1

terenzreignz (terenzreignz):

I don't know, IS IT? Check the formula I gave you.

OpenStudy (anonymous):

im so confused

OpenStudy (inkyvoyd):

@luv2chatxoxo , have you gone over the "difference quotient" yet?

terenzreignz (terenzreignz):

I don't know, IS IT? Check the formula I gave you.

terenzreignz (terenzreignz):

Okay, here's an example: Suppose I want the derivative of \(\large 5x^4\). So first, the exponent is 4, so I multiply that to the function, but I subtract 1 from the exponent... like so: \[\Large \frac{d}{dx}5x^4 = \color{red}4\cdot5x^{4\color{red}{-1}}=\color{blue}{20x^3}\]

OpenStudy (anonymous):

so is it 3x^0?

terenzreignz (terenzreignz):

Precisely ^_^ And what's x^0?

terenzreignz (terenzreignz):

Hey, you still there? You're doing great... now what is \[\Large x^0=\color{red}? \]

OpenStudy (anonymous):

1 sorry!

terenzreignz (terenzreignz):

Good. So the derivative of 3x is just 3. What about the derivative of 8? Clue: \[\Large 8 = 8x^0\] Apply the power rule again...

OpenStudy (anonymous):

8

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

would it be 8x^-1?

terenzreignz (terenzreignz):

You forgot something... \[\Large \frac{d}{dx}ax^n = \color{red}nax^{n-1}\]

OpenStudy (anonymous):

so would it just be 0?

terenzreignz (terenzreignz):

That's precisely it :P Keep that in mind. The derivative of a *constant* is *always* zero. So the derivative of 3x + 8 is just the sum of the derivatives of 3x and 8 Which are 3 and 0, respectively. so finally, what's the derivative of 3x + 8?

OpenStudy (anonymous):

3!!

OpenStudy (anonymous):

right? haha

terenzreignz (terenzreignz):

Yes. And 3 is constant, so no matter what x is, be it 4, or literally anything on the domain, the derivative is always 3.

OpenStudy (anonymous):

so we don't even have to do anything with the 4?

terenzreignz (terenzreignz):

Nope. You got lucky, the derivative of this one was a constant. Just so I know you understand, what's the derivative of \[\Large 4x^2+x\]at x = 4?

OpenStudy (anonymous):

see now would it be 8x^1 + x^-1? and do we plug in four for x?

terenzreignz (terenzreignz):

uh-oh... how did you get x^-1? The original exponent of x was 1...

OpenStudy (anonymous):

crap i mean x^0

terenzreignz (terenzreignz):

Which is...?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so would the answer be 36

terenzreignz (terenzreignz):

uhh... 8x + 1?

terenzreignz (terenzreignz):

Maybe you need to check that again...

OpenStudy (anonymous):

oh wow 33

terenzreignz (terenzreignz):

tsk tsk :P One last, just to reassure me (and yourself) that you're well on your way to mastering the power rule.

OpenStudy (anonymous):

i have another problem i can send you and i can see if i get it

OpenStudy (anonymous):

this one!! Find the derivative of f(x) = 8x2 + 11x at x = 7.

terenzreignz (terenzreignz):

Well, shoot. Go ahead and do it. Tell me your answer.

OpenStudy (anonymous):

123?

terenzreignz (terenzreignz):

Brilliant ^_^

terenzreignz (terenzreignz):

Try this one too. I have confidence in you... Derivative of \[\Large x^3 +2x^2-6x+4\] at x = 1 Go for it, tiger :P

OpenStudy (anonymous):

ah yay thank you!! quick question though

OpenStudy (anonymous):

how would you solve this one? Find the derivative of f(x) = 5/x at x = -1.

OpenStudy (anonymous):

do you just plug in -1 for x?

OpenStudy (anonymous):

so it would be -1^0?

terenzreignz (terenzreignz):

Hang on. Step-by-step. What's the derivative of 5/x ? do you know?

OpenStudy (anonymous):

noo

terenzreignz (terenzreignz):

Pity :P Let's write it this way: \[\Large \frac5x = 5x^{-1}\] does that look better? :P Power rule still applies, by the way..

OpenStudy (anonymous):

how do you get to the negative 1 though

OpenStudy (anonymous):

like x raised to it

terenzreignz (terenzreignz):

Seriously? LOL laws of exponents... \[\Large \frac1{x^n}= x^{-n}\]

OpenStudy (anonymous):

haah ok ok i remember now ive had a long morning okay so then i plug in the -1 for x correct?

terenzreignz (terenzreignz):

Not yet. First find the derivative.

OpenStudy (anonymous):

why am i getting so confsued?

terenzreignz (terenzreignz):

Because... you didn't pay attention to basics? :P

OpenStudy (anonymous):

nooo i just have to teach everything to myself lol so 5x^-1 idk what to do

terenzreignz (terenzreignz):

I said... THIS \[\Large \frac{d}{dx}ax^n = \color{blue}{nax^{n-1}}\] still applies! Even if the exponent is negative!

OpenStudy (anonymous):

ohh!!!! so 5x^-2

terenzreignz (terenzreignz):

No... remember that your n is negative! Pay attention to signs!

OpenStudy (anonymous):

so wouldnt that be -1-1? isn't that negative 2?

terenzreignz (terenzreignz):

You can't keep forgetting these things! \[\Large \frac{d}{dx}ax^n = \color{red}nax^{n-1}\]

OpenStudy (anonymous):

-5x^-2?

terenzreignz (terenzreignz):

Better. Now plug in x = -1

OpenStudy (anonymous):

so -5?

terenzreignz (terenzreignz):

mhmm ^_^

OpenStudy (anonymous):

one more :( i'm sorry i jsut need to make sure i get these Find the derivative of f(x) = -7/x at x = -3.